Using either the substitution or elimination method solve this problem. A certain aircraft can fly 1330 miles with the wind in 5 hours and travel the same distance against the wind in 7 hours. What is the speed of the wind?
can you please explain each step so I can understand where everything comes from? Thanks

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

speed of plane=P
speed of wind=W
To fly 1330 miles in 5 hours means 1330/5=266 mph
to fly 1330 miles in 7 hours means 1330/7=190 mph
The wind is added or subtracted to the plane's speed.
P+W=266; here the wind speed is added to the plane's speed.
P-W=190; subtracted.
Add the two, and the W will disappear
2P=456
divide by 2 both sides
P=228 mph.
P+W=266
W=266-228=38 mph. ANSWER
P-W=228=38=190 mph, the speed of the plane against the wind.

Answer by ikleyn(52798)

.
Using either the substitution or elimination method solve this problem. A certain aircraft can fly 1330 miles with the wind
in 5 hours and travel the same distance against the wind in 7 hours. What is the speed of the wind?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

See also the lessons
    - Wind and Current problems
    - Selected problems from the archive on a plane flying with and against the wind
in this site, where you will find similar solved problems.